The differential equation of all circles which passes through the origin and whose centre lies on y axis is

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

let the center is (o,a)  and  radius is r

so equation of circle is

x2 + (y-a)2 =r2

 differentiate

2x +2(y-a) dy/dx =0

x+ (y-a)dy/dx =0  ..............1

 

again differentiate

1+ (y-a)d2y/dx2 + dy/dx =0

put value of (y-a) from equation 1

1 -x(d2y/dx2)/(dy/dx) + dy/dx =0

(dy/dx )2 + dy/dx  - xd2y/dx2=0

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin